OpenStudy (anonymous):

Eric purchased three music CDs and a sweatshirt from a band's online store and received a 20% discount. He paid $100 for his purchase. The following week, the same store offered a 40% discount on all its products. Eric's friend Neil took advantage of the sale and bought four music CDs and two sweatshirts for $120 Assuming that the actual prices of the music CDs and the sweatshirts are unchanged, the price of a music CD is $ a. 20 b. 25 c. 30 d. 50 and the price of a sweatshirt is $ a. 20 b. 25 c. 30 d. 50

OpenStudy (anonymous):

ahhh some one has the same ??? as mine

OpenStudy (anonymous):

Let c = CD Let s= sweatshirt 3c+s=100/.8 4c+2s=120/.6 3c+s=125 4c+2s=200 Multiply the first equation by -2 -2(3c)+(-2)(s)=-2(125) -6c-2s=-250 Eliminate the s variable by adding -6c-2s=-250 to the second equation 4c+2s=200 We get the following: -2c=-50 c=25 So the regular price of the cd is $25 Plug the $25 for each cd into the first equation and solve for s. 3(25) + s = 125 s=50 So cost of a sweatshirt is $50